TileCal timing corrections are currently not in the calibration database but are available for different run ranges as data files, which have to be copied to the reconstruction directory by hand. A list of valid timing corrections can be found here.
The corresponding Tile2007.tdlas and Tile2007.tclas files can be found in the respective subdirectories here:
/afs/cern.ch/user/n/ngollub/public/timing
In addition the TileConditions package needs to be told to use calibration files with the "Tile2007" prefix in the reconstruction. This is done by adding the following two lines at the end of the RecExCommission_topOptions.py:
from TileConditions.TileConditionsConf import TileInfoLoader
TileInfoLoader.CalibFilePrefix = "Tile2007"
(The default in RecExCommission jobOptions is to use calibration files with the prefix "TileMobi", but no such files exist and thus defaults are loaded.) In order to verify which files are really loaded during running, one could also add
TileInfoLoader.OutputLevel = DEBUG
Looking for "Tile2007" in the log file should result in a lot of WARNINGs for non-existing Tile2007.* files, but the files Tile2007.tclas, Tile2007.tdlas and Tile2007.tcfib should be found and loaded. The WARNINGs about the non-existence of the other files can be safely ignored.
Dedicated page on the method, usage and results of this algorithm for cosmic muon track reconstruction from the calorimeter data only.
Towering Mapping in the Cosmic Trigger Boards
For M5 week, the cosmic boards have the following tower mapping.
For top cosmics regions the module ordering is 13, 14, 15, 16, 17, 18, 19, 20, 11, 12, 21, 22.
For bottom cosmics regions the module ordering is 45, 46, 47, 48, 49, 50, 51, 52, 43, 44, 53, 54.
During M5 the LBA cosmic boards were connected to the Muon Output.
Where 'T' indicates the tower number within the module. There are eight scalers in the cosmic board per module. The first module on the board has scaler numbers 0-8, the second module numbers 8-15, etc.
Note: The mapping was incorrect for some time. The following reflects the mapping previous to M5 (Oct 23th, 2007)
The towering mapping the cosmic trigger boards should be the same for all regions but bottom LBA and top EBC has a slightly different mapping for reasons that are still under investigation. The current mapping for top/bottom EBA and LBA and top EBC is listed as following. The mapping of LBC and EBC has not yet been verified.
For top cosmics regions the module ordering is 13, 14, 15, 16, 17, 18, 19, 20, 11, 12, 21, 22.
For bottom cosmics regions the module ordering is 45, 46, 47, 48, 49, 50, 51, 52, 43, 44, 53, 54.
Where 'T' indicates the tower number within the module. There are eight scalers in the cosmic board per module. The first module on the board has scaler numbers 0-8, the second module numbers 8-15, etc. In bottom LBA, tower 7 of LBA45 will appear on scaler 8 which should be LBA46 tower 1 for example. In the last module LBA54, tower 7 and 8 are missing.
TileCal Cosmics runs: used setups and available data
During M3 were used different triggers paired with the TileCal board trigger. TileCal trigger from EBA+LBA+LBC: 11-22 & 43-54 (8 towers in each barrel and 6 towers in each extended barrel: 8*12*4+6*12*2 = 528 towers able to give triggers) ;
A more detailed information on the runs (to be updated as analysis goes on)
10825 : 1679 events in 13 hours (2.1 ev/min);
11000 : 3912 events in 15 hours (4.2 ev/min);
Combined LAr+Tile Cosmic Runs in spring 2007
Setup (to be completed)
Easter setup Combined runs with Extended Barrel and LAr barrel and endcaps
Modules in readout:
EBA: 13, 14, 15, 16, 17, 18, 19, 20, 45, 47, 48, 49
LBA: 14, 15, 19, 20, 49, 50, 51, 52
LBC: 15, 16, 17, 18
Runs taken
Tilecal only ntuples (and log files) in /castor/cern.ch/user/m/maneira/cosmics/2007/. Reconstruction was done with
12.5.0 and tags according to TileMuonFitter#Present_software_structure.
Runs above 10825 were reconstructed with the athena setup described CosmicCommissioningReconstructionStatus.
The setup as seen in the ntuple is also given (A for LBA, C for LBC, D for EBA, E for EBC):
Run 4120:
Started at 20:18 on April 4, stopped at 8:51 on April 5, ROD problems, 1320 events;
C15-C18 D13-D49;
Run 4198:
Started at 17:08 on April 5, stopped on April 6 at 13:39, 769 events (0.6 /min);
List of combined LARG+Tile, where to find the data
The raw data files endup on the online machine pc-lar-ros-mbl-03 in directory: /tmp/ExpertWeek/data
In CASTOR they are stored in /castor/cern.ch/atlas/LargFec/BarrelP3C/
Cosmic Runs in 2005
Ntuples ready for analysis
You can get reconstructed root files for runs on /castor/cern.ch/user/t/tilebeam/cosmics/jun05 and july05. These were made using pmt ordering and not channel ordering, so the data arrays are named A13, C45 and so on. No correction was made for the double or triple-gain runs, so you should see double or triple the energy. Here's a list of runs, with info taken from http://pcata007.cern.ch:8080/mobidaq/.
We are in the process of understanding the July 2005 data. Please see/edit TileMobiCosmicAnalysisJuly05 for a list of current issues.
September 2005 Cosmics runs:
Run 3036 back-to-back trigger, variable threshold, MobiDAQ readout, 114 events, nominal PMT gain
Run 3037 back-to-back trigger, variable threshold, MobiDAQ readout, 40 K events, nominal PMT gain
Run 3038 back-to-back trigger, variable threshold, MobiDAQ readout, 31 K events, nominal PMT gain, no LBC13
Run 3039 back-to-back trigger, variable threshold, MobiDAQ readout, 40 K events, nominal PMT gain, LBC13 included in run
Run 3048 single-tower trigger, variable threshold, MobiDAQ readout, 300K events, nominal PMT gain
Run 3072 single-tower trigger, variable threshold, MobiDAQ readout, 300K events, nominal PMT gain
Run 3074 back-to-back trigger, variable threshold, MobiDAQ readout, 150 events, nominal PMT gain
Run 3225 back-to-back trigger, ROD + MobiDAQ readout, 99 events
Run 3232 back-to-back trigger, ROD + MobiDAQ readout, 3154 events
Run 3233 back-to-back trigger, ROD + MobiDAQ readout, 26374 events
Run 3244 back-to-back trigger, ROD + MobiDAQ readout, 7155 events, TDAQ errors by L1ID jumps in ROD
Run 3252 back-to-back trigger, ROD + MobiDAQ readout, 5669 events, HV off at the end of the run
Run 3265 back-to-back trigger (threshold 33 mV), ROD + MobiDAQ readout, 21243 events
Run 3306 single-tower, variable threshold, 15 Hz, ROD + MobiDAQ readout, 500K events, nominal PMT gain, HV on in modules LBA49 and LBC17 only
July 2005 Cosmics runs:
Run 2005 back-to-back trigger, variable threshold (see file below cosmics7-29.dat), 30 events, MobiDAQ
Run 2006 back-to-back trigger, variable threshold, 110 events, MobiDAQ
Run 2007 back-to-back trigger, variable threshold, 17 events, MobiDAQ, ROS event fragment
Run 2015 back-to-back trigger, variable threshold, MobiDAQ,
Run 2016 back-to-back trigger, variable threshold, long run, MobiDAQ
Run 2018 back-to-back trigger, variable threshold, MobiDAQ
Run 2021 back-to-back trigger, variable threshold, MobiDAQ, fixed L1ID problem in full event
Run 2119 back-to-back trigger, variable threshold, ROD partition
Run 2120 back-to-back trigger, variable threshold, ROD partition
Run 2140 back-to-back trigger, variable threshold, long run, ROD partition
Run 2143 back-to-back trigger, variable threshold, ROD partition
Run 2145 back-to-back trigger, variable threshold, ROD partition, correct ROD data format from here on
Run 2147 back-to-back trigger, variable threshold, ROD partition, correct ROD data format, but BCID errors observed
Run 2002 monoCIS 40k events, cap=100pF dac=200, full event number always 0
Run 2003 monoCIS 40k events, cap=5.2pF dac=100, full event number always 0
Run 2004 CIS run 35k events, full event number always 0
Run 2009 monoCIS 40k events, cap=100pF dac=200, event numbers fixed
Run 2010 monoCIS 40k events, cap=5.2pF dac=100, event numbers fixed
Run 2011 CIS run 35k events, event numbers fixed
September 2005 Pedestal runs:
Run 3030 - pedestals, 40k events, bigain, MobiDAQ readout
Run 3060 - pedestals, 40k events, bigain, MobiDAQ readout
Run 3237 - pedestals, 40k events, bigain, ROD + MobiDAQ readout, EB errors, BCID errors observed
Run 3245 - pedestals, 40k events, bigain, ROD + MobiDAQ readout, EB errors, BCID errors observed
Run 3247 - pedestals, 40k events, bigain, ROD + MobiDAQ readout
Run 3249 - pedestals, 26934 events, bigain, ROD + MobiDAQ readout
Run 3251 - pedestals, 34123 events, bigain, ROD + MobiDAQ readout, BCID errors observed
Run 3253 - pedestals, 40k events, bigain, ROD + MobiDAQ readout, EB errors, BCID errors observed
July 2005 Pedestal runs:
Run 2008 - pedestals, 10k events, bigain
June 2005 Cosmics runs:
Run 1108 is back-to-back, 35mv threshold, 45C no HV, 500 events (likely noise)
Run 1109 is single-tower triggered (50 mV, 2.0 GeV), nominal gain, about 12 hours (?)
Run 1110 Cosmics back-to-back, 84 events, threshold ?
Run 1114 is back-to-back tower triggered (threshold 38 mV, 1.5 GeV), nominal gain, 14 hours
Run 1126 is back-to-back tower triggered (threshold 38 mV, 1.5 GeV), 2*nominal gain, no 45A, 160 events
Run 1134 is back-to-back tower triggered (threshold 38 mV, 1.5 GeV), triple gain, 3115 events, almost 64 hours
June 2005 CIS runs:
Run 1102 10k events
Run 1105 monoCIS 2k events
Run 1112 35 k events, no module LBC45
Run 1154 monoCIS 1k events, trigger word, LBC45C, LBC46, LBA46
Run 1155 CIS run 35k events, trigger word, LBC45C, LBC46, LBA46
June 2005 Pedestal runs:
Run 1106 - pedestals, high gain only, 2000 events
Run 1113 - pedestal run, 35K events, no module LBC45
Summary of July 2005 Expert Week
LVPS
Three working finger power supply assemblies were used on bottom modules (LBA49, 50, 52) over several days pedestals, CIS, cosmics. The LVPS assembly for LBA51 failed and was replaced by an external supply. The reasons for the failure are being investigated.
Installation of LVPS assemblies in fingers requires special care to avoid restricting cooling.
Several LVPS assemblies passed QC at Prevessin but failed in UX15. Again, the reasons are being investigated. Future tests in Prevessin should be with long cables to include capacitance, delays, etc.
On LBA50, the +15 V MB brick tripped at 9:49PM on Sat. 30.07.05 and was restarted successfully at 10:41PM on 31.07.05. This voltage is only used for charge injection. Earlier trips like this appeared to be associated with application of HV but this one was not associated with HV activity. The -15V HV brick on LBA50 went out of regulation from 14:33 to 18:18 on 29.7.05.
Electronics noise from pedestal run (Run 2008) a. Module High gain Low gain Dead ch#[0] LVPS b. A49 1.47 0.63 - new c. A50 1.50 0.63 - new d. A51 1.17 0.48 8 (low?) old e. A52 (HV off) 1.43 0.65 - new w cores f. C17 1.23 0.54 25 old g. C18 1.24 0.51 - old h. C19 1.14 0.51 - old i. C20 1.15 0.51 - old
ROD
One ROD successfully tested to work in the LBC partition together with a TBM. Stand-alone mode with injected fake data no errors. Successfully integrated ROD in the online software 01-02-00. Combined partition to start both MobiDAQ and ROD segments. Many changes applied to the DSP code in order to get readable output data for the offline.
Could not upgrade monitoring applications on time for cosmic runs. Linkage errors with one of the fibres from the FE (FragId 0x130). Combined event building problems (smp kernel or network failure + wrong BCID because of fragment loss in fibre 0x130). Data from the last run corrupted since HV trip.
Front-End Electronics
C17 Pulse shape of channel 26 (25 counting from 0) low gain problem: all zeroes, both for monoCIS and pedestals. High gain looks fine. Likely a hardware problem in a digitizer?
HV
Constant overcurrent tripped state on 1 module (52A) Wild current fluctuations on 1 module (50A) o Verified with 2 different LV supplies (old+new) One trip at end (49A) after 3 days
DAQ
bug found in DAQ software, resulting in event number always zero for some of the data.
HLT
first use of event filter in PC
Cooling
Heater failed in cooling unit on C-side, replaced with temporary heater, part on order (end August). System ran stably throughout the data taking.
DCS
Difficult to recover from power cut / reset DCS to racks Need better on-line monitoring of LV, HV a. LVPS assembly monitoring is good but not integrated with external LV supplies, HV, Temperatures, or run data base b. Work needed on sense-line readout for MB voltages
Offline
Need to store HV values read from drawer during run into database
Example event displays
See attached files at bottom.
Cosmic Run 7814 event 475. Example of a shower candidate: